# Analysis of a class of Recursive Distributional Equations Including the Resistance of the Series-parallel graph

### v1

@@ A.2. Uniqueness for Nondecreasing Data
If $F_{\text{in}}$ is merely absolutely continuous with $\partial_{x}F_{\text{in}}\in L^{1}(\mathbb{R})$, nonetheless $\partial_{x}F_{\text{in}}$ can be approximated in $L^{1}(\mathbb{R})$ by smooth, nowhere-vanishing functions. The $L^{1}$ contraction property of the PME then implies that the corresponding solutions converge in $L^{\infty}([0,\infty);L^{1}(\mathbb{R}))$, which implies uniform convergence of the antiderivatives. Thus, since the viscosity solution property is stable under uniform convergence, even if $\partial_{x}F_{\text{in}}$ is merely in $L^{1}(\mathbb{R})$, the unique viscosity solution of (57) in this case is still given as the antiderivative of $\rho$.
Finally, it remains to consider the case when $\partial_{x}F_{\text{in}}$ is only a finite positive measure (so that $F_{\text{in}}$ may have discontinuities). Let $\{F^{(N)}_{\text{in}}\}\subseteq CDF(\overline{\mathbb{R}})$ be smooth functions such that $F^{(N)}_{\text{in}}\leq F^{(N+1)}_{\text{in}}<F_{\text{in}}$ for each $N$, $F^{(N)}_{\text{in}}\to F_{\text{in}}$ at points of continuity as $N\to\infty$. If $\rho_{N}$ is the corresponding solution of the PME (64) and $F_{N}$, its antiderivative, then $F_{N}\leq\overline{F}$ for each $N$ since $F_{N}$ is a viscosity subsolution of (57).
At the same time, some subsequence of $\{\rho_{N}\}$ converges locally uniformly in $\mathbb{R}\times(0,\infty)$ as $N\to\infty$ to a continuous, nonnegative distributional solution $\rho$ of (64) with initial trace equal to the measure $\partial_{x}F_{\text{in}}$; this follows, for instance, from the fact that, in this one-dimensional setting, the pressure $\{\rho_{N}^{m-1}\}$ is both uniformly bounded and uniformly Lipschitz in $\mathbb{R}\times[\tau,+\infty)$ for any $\tau>0$, independently of $N$ (see [21, Chapter 14]). A classical result of Dahlberg and Kenig [14] implies that $\rho$ is unique, hence the entire sequence $\{\rho_{N}\}$ converges to $\rho$. At the level of antiderivatives, this implies $F_{N}\to F$ as $N\to\infty$, where $F$ is given by (64) and is a discontinuous viscosity solution of (57). By comparison, $\overline{F}\leq F$, hence actually $F=\overline{F}$.
Approximating $F_{\text{in}}$ from above, one similarly finds $F=\underline{F}$. ∎
@@ A.3. Viscosity Solutions in $CDF(\overline{\mathbb{R}})$
Appendix B Scaling Limits of Monotone Semigroups in $CDF(\overline{\mathbb{R}})$
This appendix proves Theorem 3, the extension of the convergence framework of [6] to approximations of PDE’s posed in $CDF(\overline{\mathbb{R}})$. As described already briefly in the introduction, technical issues arise stemming from the fact that $CDF(\overline{\mathbb{R}})$ is not a vector space, and the smooth test functions used in the definition of viscosity solution are not necessarily nondecreasing. These issues are not fatal and, indeed, similar ones have been overcome in other settings, such as geometric flows (see [9]).
@@ Appendix B Scaling Limits of Monotone Semigroups in $CDF(\overline{\mathbb{R}})$
Let $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a continuous function that is nondecreasing in the second argument. An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ satisfies $\partial_{t}F\leq\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)$ in the viscosity sense if and only if it has the following property: If $\varphi$ is a smooth test function that touches $F$ from above at some point $(x_{0},t_{0})$ and
@@
The fact that this weaker definition of viscosity subsolution is equivalent to the usual one can be proved exactly as in [8, Proposition 2.2] (see also [20, Proof of Theorem 1.2]). ∎
@@
A corresponding version of Lemma 3 holds for supersolutions, which can be deduced from the transformation G(x,t)=1F(x,t)G(x,t)=1-F(-x,t).
@@ B.1. Proof of Theorem 3
Since $F_{\epsilon}$ is discontinuous in time, it will clarify the picture somewhat to observe that the statement that $F_{\epsilon}\to F$ locally uniformly in $\mathbb{R}\times[0,\infty)$ is equivalent to the following: For each $(x,t)\in\mathbb{R}\times[0,\infty)$,
(65) $\lim_{\delta\downarrow 0}\sup\left\{|F^{\epsilon}_{n}(y)-F(x,t)|\,\mid\,|y-x|+|\epsilon^{\alpha}n-t|\leq\delta\right\}=0.$
This is the precise statement that will be proved below.
@@
Since the operators $(T_{\epsilon})_{\epsilon>0}$ are monotone and nondecreasing continuous functions can be approximated locally uniformly from above and below by smooth functions, it suffices to consider the case when $F^{\epsilon}_{0}=F_{\text{in}}$ for each $\epsilon>0$ and $F_{\text{in}}$ has a bounded, uniformly continuous second derivative. This will be assumed for the remainder of the proof.
Throughout what follows, define the constants $c_{-},c_{+}\in[0,1]$ by
(66) $c_{-}=F_{\text{in}}(-\infty)=F_{\epsilon}(+\infty,t),\quad c_{+}=F_{\text{in}}(+\infty)=F_{\epsilon}(-\infty,t),$
where property (iii) implies that the identities above hold independently of $t>0$.
@@
Step 1: Lipschitz Estimate in Time. Since $F_{\text{in}}$ is currently assumed to have a bounded, uniformly continuous second derivative, by the consistency condition (iv),
$\epsilon^{-\alpha}\mathscr{L}_{\epsilon}F_{0}^{\epsilon}=\epsilon^{-\alpha}\mathscr{L}_{\epsilon}F_{\text{in}}=\mathcal{F}(\partial_{x}F_{\text{in}},\partial^{2}_{x}F_{\text{in}})+R_{\epsilon}$
@@
where $R_{\epsilon}(x)\to 0$ uniformly with respect to $x$ as $\epsilon\to 0$. Thus, since $\mathcal{F}$ is bounded on compact sets by continuity, there is a constant $C_{0}>0$ such that, for sufficiently small $\epsilon$,
$\|T_{\epsilon}F_{0}^{\epsilon}-F_{0}^{\epsilon}\|_{\text{sup}}\leq C_{0}\epsilon^{\alpha}\quad\text{in}\,\,\mathbb{R}.$
In Proposition 12 below, the assumptions (i)-(iii) are shown to imply the following contractivity property of $T_{\epsilon}$:
$\|T_{\epsilon}F-T_{\epsilon}G\|_{\text{sup}}\leq\|F-G\|_{\text{sup}}.$
Thus, the estimate of $T_{\epsilon}F_{0}^{\epsilon}-F_{0}^{\epsilon}$ may be iterated to find
$\|F^{\epsilon}_{n+1}-F_{n}^{\epsilon}\|_{\text{sup}}=\|T_{\epsilon}F^{\epsilon}_{n}-F_{n}^{\epsilon}\|_{\text{sup}}\leq C_{0}\epsilon^{\alpha}\quad\text{for each}\,\,n\in\mathbb{N}_{0}.$
Note, in particular, that $|F^{\epsilon}_{n}(x)-F_{\text{in}}(x)|\leq C_{0}\epsilon^{\alpha}n$ uniformly in $(x,n)$ as $\epsilon\to 0$.
@@
Step 3: Convergence. The proof uses the classical half-relaxed limit method. Let $F_{\star}$ and $F^{\star}$ be the half-relaxed limits of $\{F^{\epsilon}_{n}\}$ defined by
$F_{\star}(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{n}^{\epsilon}(y)\,\mid\,|x-y|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\},$
$F^{\star}(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{n}^{\epsilon}(y)\,\mid\,|x-y|+|\epsilon^{\alpha}n-t|+\epsilon\leq\delta\right\},$
@@
To prove the local convergence statement (65), it suffices to show that $F^{\star}=F_{\star}=F$ (see [13, Remark 6.4] or [7, Lemma 6.2]). To do this, since $F_{\star}\leq F^{\star}$ is immediate, it is only necessary to show that $F^{\star}\leq F\leq F_{\star}$. This will follow from the comparison principle, Theorem 4, as soon as it is established that $F^{\star}$ and $F_{\star}$ define respectively a viscosity sub- and a viscosity supersolution of the initial value problem (57).
@@
Step 4: $F^{\star}$ is a Subsolution. In view of what was proved in Step 1, it is clear that $F^{\star}(\cdot,0)=F_{\text{in}}$ in the sense of (58). Therefore, it only remains to establish that $\partial_{t}F^{\star}-\mathcal{F}(\partial_{x}F^{\star},\partial_{x}^{2}F^{\star})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$.
@@
In order to avoid boundary issues that can arise if $F^{\star}$ is somewhere equal to one, it will be convenient to pull $F^{\star}$ down: To see that $F^{\star}$ is a subsolution, it suffices to prove that, for any $\delta\in(0,c_{+}-c_{-})$, the function $G^{\star}:\mathbb{R}\times[0,\infty)\to[c_{-},c_{+}-\delta]$ defined by $G^{\star}=\max\{c_{-},F^{\star}-\delta\}$ satisfies
$\partial_{t}G^{\star}\leq\mathcal{F}(\partial_{x}G^{\star},\partial_{x}^{2}G^{\star}).$
@@
For the rest of this step, fix $\delta$ and consider $G^{\star}$. Note also that if $\{G^{\epsilon}_{n}\}$ is the sequence defined for $\epsilon>0$ by $G^{\epsilon}_{n}=\max\{c_{-},F^{\epsilon}_{n}-\delta\}$, then Lemma 4 below implies that
(67) $G^{\epsilon}_{n+1}\leq T_{\epsilon}G^{\epsilon}_{n}\quad\text{pointwise for each}\,\,n\in\mathbb{N}_{0}.$

### v2

@@ A.2. Uniqueness for Nondecreasing Data
If $F_{\text{in}}$ is absolutely continuous with $\partial_{x}F_{\text{in}}\in L^{1}(\mathbb{R})$, then it can be approximated by smooth, strictly increasing functions $\{F_{\text{in}}^{(N)}\}$, with $\{\partial_{x}F_{\text{in}}^{(N)}\}$ converging in $L^{1}$, and the corresponding solutions $\{\rho^{(N)}\}$ of the PME then converge in $L^{\infty}([0,+\infty);L^{1}(\mathbb{R}))$ by the $L^{1}$ contractivity property of the equation (see [24, Chapter 9]) to the solution $\rho$ started from $\partial_{x}F_{\text{in}}$. This convergence implies uniform convergence of the antiderivatives: Since the uniform limit of viscosity solutions is a viscosity solution, once again (72) provides the viscosity solution in this case. Note also that item (i) remains true and the bounds (76) are preserved in the limit.
Finally, if $\partial_{x}F_{\text{in}}$ is merely a finite measure, consider sequences of smooth functions $\{F^{(N),+}_{\text{in}}\}$ and $\{F^{(N),-}_{\text{in}}\}$ as in Lemma 1, chosen in such a way that $F^{(N),\pm}_{\text{in}}(\bar{x})\to F_{\text{in}}(\bar{x})$ as $N\to+\infty$ for each $\bar{x}\in\{-\infty,+\infty\}$. By the bounds (76), the corresponding solutions of the PME $\{\rho^{(N),+}\}$ and $\{\rho^{(N),-}\}$ started from the data $\{\partial_{x}F^{(N),+}_{\text{in}}\}$ and $\{\partial_{x}F^{(N),-}\}$ are uniformly bounded and equicontinuous in $\mathbb{R}\times(\tau,+\infty)$ for any $\tau>0$. Therefore, the antiderivatives $\{F^{(N),+}\}$ and $\{F^{(N),-}\}$ become equicontinuous after a positive amount of time. Since they are, respectively, monotone decreasing and monotone increasing in $N$, the limits $F^{+}=\lim_{N\to\infty}F^{(N),+}$ and $F^{-}=\lim_{N\to\infty}F^{(N),-}$ exist. Since $F^{(N),-}\leq F^{-}\leq F^{+}\leq F^{(N),+}$ holds in $\mathbb{R}\times[0,+\infty)$ for any $N$, $F^{-}(\cdot,t)$ and $F^{+}(\cdot,t)$ both converge vaguely to $F_{\text{in}}$ as $t\downarrow 0$. By Arzelà-Ascoli, the derivatives $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ are continuous distributional solutions of the PME in $\mathbb{R}\times(0,+\infty)$, and the vague convergence of $F^{-}$ and $F^{+}$ at time $t=0$ implies that the initial trace of $\partial_{x}F^{-}$ and $\partial_{x}F^{+}$ equals $\partial_{x}F_{\text{in}}$. Therefore, by Dahlberg-Kenig, $\partial_{x}F^{-}=\partial_{x}F^{+}$ in $\mathbb{R}\times(0,+\infty)$, which implies the difference $F^{+}-F^{-}$ depends only on time.
Since item (i) holds for regular solutions, and, in the construction above, $F^{(N),\pm}_{\text{in}}(\bar{x})\to F_{\text{in}}(\bar{x})$ for each $\bar{x}\in\{-\infty,+\infty\}$, it follows that $F^{+}(\bar{x},t)=F^{-}(\bar{x},t)=F_{\text{in}}(\bar{x})$ also holds for any $t$. Therefore, $F^{+}=F^{-}$, and so (72) follows by Lemma 1 and item (i) holds.
Item (ii) follows as in the previous proposition.∎
Case $Q\neq 0$. Next, using only the comparison principle, uniqueness in the case $Q=0$ is upgraded to the case when $Q$ is nonzero.
Proposition 16.
The identity $\overline{F}=\underline{F}$ in $\mathbb{R}\times(0,\infty)$ still holds for the initial-value problem (61) with arbitrary data in $CDF(\overline{\mathbb{R}})$ if $\mathcal{F}$ satisfies the assumptions of Propositions 14 or 15 and $Q$ satisfies (63) and (64).
Proof.
Fix $F_{\text{in}}\in CDF(\overline{\mathbb{R}})$. If $F_{\text{in}}$ is continuous, then Proposition 12 implies there is a unique viscosity solution $F$ of (61) and $\overline{F}=\underline{F}=F$. Otherwise, if it has jumps, fix sequences $\{F_{\text{in}}^{(N),+}\}$ and $\{F_{\text{in}}^{(N),-}\}$ of continuous functions in $CDF(\overline{\mathbb{R}})$ satisfying (67) and (68) as in Lemma 1 and, in addition,
(77) $\lim_{N\to\infty}F_{\text{in}}^{(N),-}(\pm\infty)=\lim_{N\to\infty}F_{\text{in}}^{(N),+}(\pm\infty)=F_{\text{in}}(\pm\infty).$
For each $N$, let $F^{(N),+}$ and $F^{(N),-}$ denote the unique viscosity solutions of (61) with initial data $F_{\text{in}}^{(N),+}$ and $F_{\text{in}}^{(N),-}$, respectively.
By construction and the comparison principle, the bounds $F^{(N),-}\leq\underline{F}\leq\overline{F}\leq F^{(N),+}$ hold pointwise in $\mathbb{R}\times[0,+\infty)$. Therefore, it only remains to show that, for any $t>0$,
(78) $\lim_{N\to\infty}\sup\left\{F^{(N),+}(x,t)-F^{(N),-}(x,t)\,\mid\,x\in\mathbb{R}\right\}=0.$
To this end, since $Q$ is bounded in $[0,1]$, notice that the modified functions $\tilde{F}^{(N),+}$ and $\tilde{F}^{(N),-}$ given by
$\tilde{F}^{(N),\pm}(x,t)=F^{(N),\pm}(x,t)\mp C(Q)t,\quad\text{where}\quad C(Q)=\sup\left\{|Q(q)|\,\mid\,0\leq q\leq 1\right\},$
are, respectively, sub- and supersolutions of the PDE $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0$ with $Q$ removed. Therefore, if $G^{(N),+}$ and $G^{(N),-}$ are the viscosity solutions of this PDE corresponding to initial data $F^{(N),+}_{\text{in}}$ and $F^{(N),-}_{\text{in}}$, respectively, then the comparison principle implies
(79) $F^{(N),+}-C(Q)t=\tilde{F}^{(N),+}\leq G^{(N),+}\quad\text{and}\quad G^{(N),-}\leq\tilde{F}^{(N),-}=F^{(N),-}+C(Q)t,$
where all the inequalities above hold pointwise in $\mathbb{R}\times[0,+\infty)$. On the other hand, by Propositions 14 or 15, if $G$ is the discontinuous viscosity solution of the PDE $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)=0$ with initial datum $F_{\text{in}}$, then, for any $\delta>0$, the functions $G^{(N),+}(\cdot,\delta)$ and $G^{(N),-}(\cdot,\delta)$ converge pointwise in $\mathbb{R}$ to $G(\cdot,\delta)$. Since the constants $G^{(N),+}(\pm\infty,\delta)$ and $G^{(N),-}(\pm\infty,\delta)$ also converge to $G(\pm\infty,\delta)$ by (77) and item (i) in the aforementioned propositions, Dini’s Theorem (applied to the compact metric space $\mathbb{R}\cup\{-\infty,+\infty\}$) implies that this convergence is actually uniform:
$\lim_{N\to\infty}\sup\left\{G^{(N),+}(x,\delta)-G^{(N),-}(x,\delta)\,\mid\,x\in\mathbb{R}\right\}=0.$
Therefore, by (79),
(80) $\lim_{N\to\infty}\sup\left\{F^{(N),+}(x,\delta)-F^{(N),-}(x,\delta)\,\mid\,x\in\mathbb{R}\right\}\leq 2C(Q)\delta.$
Finally, moving back to the original problem (i.e., with the reaction term $Q$), the estimate (66) in Theorem 4 implies that, for any $t>\delta$,
$\limsup_{N\to\infty}\sup\left\{F^{(N),+}(x,t)-F^{(N),-}(x,t)\,\mid\,x\in\mathbb{R}\right\}\leq 2C(B)\delta e^{L(t-\delta)}.$
After sending $\delta\downarrow 0$, this becomes (78), as desired. ∎
@@ A.3. Viscosity Solutions in $CDF(\overline{\mathbb{R}})$
Appendix B Scaling Limits of Some Monotone Semigroups in $CDF(\overline{\mathbb{R}})$
This appendix proves Theorem 3, an extension of the convergence framework of [7] to approximations of PDE’s posed in $CDF(\overline{\mathbb{R}})$. As described already briefly in the introduction, technical issues arise stemming from the fact that $CDF(\overline{\mathbb{R}})$ is not a vector space, and the smooth test functions used in the definition of viscosity solution are not necessarily nondecreasing. These issues are not fatal and, indeed, similar ones have been overcome in other settings, such as geometric flows (see [9]).
@@ Appendix B Scaling Limits of Some Monotone Semigroups in $CDF(\overline{\mathbb{R}})$
Let $\mathcal{F}:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a continuous function that is nondecreasing in the second argument and $Q:\mathbb{R}\to\mathbb{R}$ be continuous. An upper semicontinuous function $F:\mathbb{R}\times(0,\infty)\to\mathbb{R}$ satisfies $\partial_{t}F-\mathcal{F}(\partial_{x}F,\partial^{2}_{x}F)+Q(F)\leq 0$ in the viscosity sense if and only if it has the following property: If $\varphi$ is a smooth test function that touches $F$ from above at some point $(x_{0},t_{0})$ and
@@
The fact that this weaker definition of viscosity subsolution is equivalent to the usual one can be proved exactly as in [8, Proposition 2.2] (see also [23, Proof of Theorem 1.2]). ∎
@@
A corresponding version of Lemma 4 holds for supersolutions, as can be readily checked using the transformation G(x,t)=1F(x,t)G(x,t)=1-F(-x,t).
@@ B.1. Proof of Theorem 3
It is technically convenient to proceed by proving the following statement: For each $(x,t)\in\mathbb{R}\times[0,\infty)$,
(81) $\lim_{\delta\downarrow 0}\sup\left\{|F^{(N)}_{n}(y)-F(x,t)|\,\mid\,|\frac{y}{\delta^{(N)}}-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\}=0.$
Notice that this implies that $F_{N}\to F$ locally uniformly in $\mathbb{R}\times[0,+\infty)$ as $N\to+\infty$.
@@
Since the operators $\{T^{(N)}\}$ are monotone, and nondecreasing continuous functions can be approximated locally uniformly from above and below by smooth functions, it suffices to consider the case when equality holds $F^{(N)}_{0}=F_{\text{in}}$ for each $N$ and $F_{\text{in}}$ has bounded, uniformly continuous first and second derivatives. This will be assumed for the remainder of the proof.
Throughout what follows, define functions $c_{-},c_{+}:[0,\infty)\to[0,1]$ by
(82) $\frac{dc_{\pm}}{dt}=-Q(c_{\pm}),\quad c_{+}(0)=F_{\text{in}}(-\infty),\quad c_{+}(0)=F_{\text{in}}(+\infty).$
Assumption (vi) ensures that these ODE’s are wellposed. For each $N$, let $\{c_{-}^{(N)}(n)\}$ and $\{c_{+}^{(N)}(n)\}$ be the sequences of numbers in $[0,1]$ defined analogously by
$c_{\pm}^{(N)}(n)=T^{(N)}c_{\pm}(n-1)=c_{\pm}^{(N)}(n-1)-Q^{(N)}(c_{\pm}^{(N)}(n-1)),$
$c_{+}^{(N)}(0)=c_{+}(0),\quad c_{-}^{(N)}(0)=c_{-}(0).$
(Assumptions (i) and (iv) imply that $T^{(N)}$ maps constant functions to constant functions, and $\mathcal{L}^{(N)}$ vanishes on constants, as explained in the discussion in Section 5.1.) By assumption (vi) and standard ODE arguments, for any $T>0$,
(83) $\lim_{\delta\downarrow 0}\sup\left\{|c_{\pm}^{(N)}(n)-c_{\pm}(t)|\,\mid\,|N^{-1}n-t|+N^{-1}\leq\delta,\,\,t\in[0,T]\right\}=0.$
Further, by assumption (iv), for any $N$,
(84) $c_{-}^{(N)}(n)=F^{(N)}_{n}(-\infty)\leq F^{(N)}_{n}\leq F^{(N)}_{n}(+\infty)=c_{+}^{(N)}(n)\quad\text{for each}\,\,n\geq 0.$
@@
Step 1: Lipschitz Estimate in Time. Since $F_{\text{in}}$ is currently assumed to have bounded, uniformly continuous first and second derivatives, by the consistency condition (v),
$N\mathscr{L}^{(N)}F_{0}^{(N)}=N\mathscr{L}^{(N)}F_{\text{in}}=\mathcal{F}(\partial_{x}F_{\text{in}},\partial^{2}_{x}F_{\text{in}})+R^{(N)}$
@@
where $R^{(N)}\to 0$ uniformly as $N\to+\infty$. Thus, since $\mathcal{F}$ is bounded on compact sets by continuity and the reaction term $Q^{(N)}$ is of order $N^{-1}$ by assumption (vi), there is a constant $C_{0}>0$ such that, for sufficiently large $N$,
$\|T^{(N)}F_{0}^{(N)}-F_{0}^{(N)}\|_{\text{sup}}\leq C_{0}N^{-1}.$
By Proposition 17 below, the assumptions (i)-(iv) and (vi) imply the following contractivity property of $T^{(N)}$:
$\|T^{(N)}F-T^{(N)}G\|_{\text{sup}}\leq(1+LN^{-1})\|F-G\|_{\text{sup}}.$
Thus, the estimate of $F^{(N)}_{1}-F_{0}^{(N)}=T^{(N)}F_{0}^{(N)}-F_{0}^{(N)}$ can be iterated to find
$\|F^{(N)}_{n+1}-F_{n}^{(N)}\|_{\text{sup}}=\|T^{(N)}F^{(N)}_{n}-T^{(N)}F_{n-1}^{(N)}\|_{\text{sup}}\leq C_{0}N^{-1}(1+N^{-1}L)^{n}\quad\text{for each}\,\,n\in\mathbb{N}.$
Summing this estimate in $n$, one obtains a constant $C_{0}^{\prime\prime}>0$ such that that $|F^{(N)}_{n}(x)-F_{\text{in}}(x)|\leq C_{0}^{\prime\prime}N^{-1}(1+N^{-1}L)^{n}$ uniformly in $(x,n)$ as $N\to+\infty$. In view of the definition of $F_{N}$, this implies that, for each $T>0$,
$\sup\left\{|F_{N}(x,t)-F_{\text{in}}(x)|\,\mid\,(x,t)\in\mathbb{R}\times[0,T]\right\}\leq C_{0}^{\prime\prime}N^{-1}(1+LN^{-1})^{[NT]+1}.$
@@
Step 2: Convergence. The remainder of the proof uses the classical half-relaxed limit method. Let $F_{\star}$ and $F^{\star}$ be the half-relaxed limits of $\{F^{(N)}_{n}\}$ defined by
$F_{\star}(x,t)$ $=\lim_{\delta\downarrow 0}\inf\left\{F_{n}^{(N)}(y)\,\mid\,|\delta^{(N)}y-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\},$
$F^{\star}(x,t)$ $=\lim_{\delta\downarrow 0}\sup\left\{F_{n}^{(N)}(y)\,\mid\,|\delta^{(N)}y-x|+|N^{-1}n-t|+N^{-1}\leq\delta\right\},$
@@
As is well-known, to prove the local convergence statement (81), it suffices to show that $F^{\star}=F_{\star}=F$ (see [15, Remark 6.4] or [10, Lemma 6.2]). To do this, since the bound $F_{\star}\leq F^{\star}$ follows immediately from the definitions, it is only necessary to show that $F^{\star}\leq F\leq F_{\star}$. This will follow from the comparison principle, Theorem 4, as soon as it is established that $F^{\star}$ and $F_{\star}$ define respectively a viscosity sub- and a viscosity supersolution of the initial value problem (61).
@@
Step 3: $F^{\star}$ is a Subsolution. In view of what was proved in Step 1 and the uniform continuity of $F_{\text{in}}$, it is clear that $F^{\star}$ coincides with $F_{\text{in}}$ in the sense of (65). Therefore, to see that $F^{\star}$ is a viscosity subsolution of (61), it only remains to establish that $\partial_{t}F^{\star}-\mathcal{F}(\partial_{x}F^{\star},\partial_{x}^{2}F^{\star})+Q(F^{\star})\leq 0$ in the viscosity sense in $\mathbb{R}\times(0,\infty)$.
@@
In order to avoid boundary issues that can arise if $F^{\star}$ is somewhere equal to one, it will be convenient to pull $F^{\star}$ down: To see that $F^{\star}$ is a subsolution, it suffices to prove that, for any $\delta\in(0,1)$, the function $G^{\star}:\mathbb{R}\times[0,\infty)\to[0,1-\delta]$ defined by $G^{\star}(x,t)=\max\{F^{\star}(x,t)-\delta e^{Lt},0\}$ satisfies
$\partial_{t}G^{\star}-\mathcal{F}(\partial_{x}G^{\star},\partial_{x}^{2}G^{\star})+Q(G^{\star})\leq 0\quad\text{in the viscosity sense in}\,\,\mathbb{R}\times(0,+\infty).$
@@
(By standard arguments, the subsolution property of $F^{\star}$ can then be recovered upon sending $\delta\downarrow 0$.) For the rest of this step, fix $\delta$ and consider $G^{\star}$. Note also that if $\{G^{(N)}_{n}\}$ is the sequence defined for a given $N$ by $G^{(N)}_{n}=\max\{F^{(N)}_{n}-\delta(1+LN^{-1})^{n},0\}$, then Lemma 5 below implies that
(85) $G^{(N)}_{n}\leq T^{(N)}G^{(N)}_{n-1}\quad\text{pointwise for each}\,\,n\in\mathbb{N}.$